# Nonlocal van der Waals density functional made faster

### v1

@@ I Introduction
Density functional theory is where the worlds of mathematics, physics, and chemistry meet in the most fruitful way, the rigorous theorems are followed by models ranging from approximations and interpolations between known limits down to empirical parametrizations of all kinds, to come up with a computational tool that helps the chemist to understand the structure and behavior of molecular systems and even to design new ones. Starting from the uniform electron gas Hohenberg and Kohn (1964); Kohn and Sham (1965), the generalized-gradient approximations Langreth and Mehl (1983); Becke (1986); Perdew and Wang (1986); Perdew, Burke, and Ernzerhof (1996) have brought the accuracy to a chemically meaningful level; a mixing of exact exchange Becke (1993a, b) or an even better long-range exchange correction Iikura et al. (2001); Heyd, Scuseria, and Ernzerhof (2003); Henderson, Janesko, and Scuseria (2008); Laikov (2019a) is the next step in reaching near chemical accuracy. What is still missing is the small but sometimes important dispersion (also known as van der Waals) energy contribution for which density functionals evolved Rydberg et al. (2000, 2003); Dion et al. (2004) in the form of a six-dimensional integral — a line of further simplifications Vydrov, Wu, and Van Voorhis (2008); Vydrov and Van Voorhis (2009, 2010a) (notwithstanding formal controversies Langreth and Lundqvist (2010); Vydrov and Van Voorhis (2010b)) has led to an elegant analytical expression for its integrand — but even then its direct numerical evaluation is rather time-consuming. In plane-wave methods a fast Fourier transform techniques Román-Pérez and Soler (2009); Sabatini, Gorni, and de Gironcoli (2013) can help, but here we want to deal with isolated molecular systems without any further atoms-in-molecules Gräfenstein and Cremer (2009); Steinmann and Corminboeuf (2010) kind of approximation.
@@ II Theory
of the interelectronic distance $|{\rm\bf r}_{1}-{\rm\bf r}_{2}|$ and the electron density $\rho({\rm\bf r})$ and its gradient norm $\gamma({\rm\bf r})\equiv\bigl|\nabla\rho({\rm\bf r})\bigr|$ at both points ${\rm\bf r}_{1}$ and ${\rm\bf r}_{2}$. Though the function $f(r,\rho_{1},\gamma_{1},\rho_{2},\gamma_{2})$ may have a simple enough form, the integral has to be computed by six-dimensional numerical integration. We want to find a fast approximation $\tilde{E}_{6}\approx E_{6}$ as a sum of a few integrals
@@
$B\equiv\tfrac{1}{2}3^{2/3}\pi^{5/6}b$, with two adjustable parameters $C$ and $b$, and the function of the squared distance is
@@
Our first approximation assumes $\omega_{1}\approx\omega_{2}$, helping to simplify the function down to three variables
@@
and the new function of the distance
$v(s,\mu_{1},\mu_{2})$
$=$ $\frac{1}{(s+\mu_{1})(s+\mu_{2})\bigl(s+\tfrac{1}{2}(\mu_{1}+\mu_{2})\bigr)}.$
@@
Our next approximation deals with the function (II)
@@
$\tilde{v}(s,\mu,\nu,\beta)=\frac{1}{(s+\beta)^{3}}+\sum_{m=4}^{n}\frac{c_{m}(\mu,\nu,\beta)}{(s+\beta)^{m}}.$ (16)
At first we hoped the simplest one with only the first term might be ideal: it has the $r^{-6}$-tail the dispersion functionals are all about, and only one spatial distribution (11) is needed. This hope was forlorn after the tests on noble-gas homo- and heterodimers: the value of $\beta$ had to vary from system to system by more than twofold to reproduce well enough the potential curves. Thus the length-scale distribution (12) should be somehow accounted for.
@@
What also matters is the sum rule, from the diagonal ($\mu=\nu$) case
@@
For $\mu\neq\nu$ we take the weighted arithmetic and geometric mean
@@
Even though $\mu({\rm\bf r})\approx\beta$ almost everywhere, it can be too big or too small somewhere, for greater numerical stability we prefer to use safer smoothed values
@@
$\theta(x,c,a)=c+ac\tanh\left(\frac{x-c}{ac}\right).$ (29)
@@
are computed using the points $\{{\rm\bf r}_{k}\}$ of a three-dimensional integration grid with weights $\{w_{k}\}$, fast multipole methods Rokhlin (1985) should help here.
@@ III Calculations
Homo- and heteroatomic noble-gas dimers are our favorite model system for testing dispersion functionals. We use our long-range-corrected version Laikov (2019a) of the PBE Perdew, Burke, and Ernzerhof (1996) exchange functional, our two-component scalar-relativistic approximation Laikov (2019b), and our L2a_3 basis set Laikov (2019c), and calculate the values shown in Table 1.

### v2

@@ I Introduction
Density functional theory is where the worlds of mathematics, physics, and chemistry meet in the most fruitful way, the rigorous theorems are followed by models ranging from approximations and interpolations between known limits down to empirical parametrizations of all kinds, to come up with a computational tool that helps the chemist to understand the structure and behavior of molecular systems and even to design new ones. Starting from the uniform electron gas Hohenberg and Kohn (1964); Kohn and Sham (1965), the generalized-gradient approximations Langreth and Mehl (1983); Becke (1986); Perdew and Wang (1986); Perdew et al. (1996) have brought the accuracy to a chemically meaningful level; a mixing of exact exchange Becke (1993a, b) or an even better long-range exchange correction Iikura et al. (2001); Heyd et al. (2003); Henderson et al. (2008); Laikov (2019c) is the next step in reaching near chemical accuracy. What is still missing is the small but sometimes important dispersion (also known as van der Waals) energy contribution for which density functionals evolved Rydberg et al. (2000, 2003); Dion et al. (2004) in the form of a six-dimensional integral — a line of further simplifications Vydrov et al. (2008); Vydrov and Van Voorhis (2009, 2010a) (notwithstanding formal controversies Langreth and Lundqvist (2010); Vydrov and Van Voorhis (2010b)) has led to an elegant analytical expression for its integrand — but even then its direct numerical evaluation is rather time-consuming. In plane-wave methods a fast Fourier transform techniques Román-Pérez and Soler (2009); Sabatini et al. (2013) can help, but here we want to deal with isolated molecular systems without any further atoms-in-molecules Gräfenstein and Cremer (2009); Steinmann and Corminboeuf (2010) kind of approximation.
@@ II Theory
of the interelectronic distance $|{\rm\bf r}_{1}-{\rm\bf r}_{2}|$ and the electron density $\rho({\rm\bf r})$ and its gradient norm $\gamma({\rm\bf r})\equiv\bigl|\nabla\rho({\rm\bf r})\bigr|$ at both points ${\rm\bf r}_{1}$ and ${\rm\bf r}_{2}$. A uniform-electron-gas compensation term
$E_{-6}=-\int\int f\bigl(|{\rm\bf r}_{1}-{\rm\bf r}_{2}|,\rho({\rm\bf r}_{1}),0,\rho({\rm\bf r}_{1}),0\bigr)\,\mathrm{d}^{3}{\rm\bf r}_{1}\,\mathrm{d}^{3}{\rm\bf r}_{2}$ (3)
is also added to the energy.
Though the function $f(r,\rho_{1},\gamma_{1},\rho_{2},\gamma_{2})$ may have a simple enough form, the integral has to be computed by six-dimensional numerical integration. We want to find a fast approximation $\tilde{E}_{6}\approx E_{6}$ as a sum of a few integrals
@@
$B\equiv\tfrac{1}{2}3^{2/3}\pi^{5/6}b$, with two adjustable parameters $C$ and $b$, and the function of the squared distance $s$ is
@@
The compensation term (3) is then simply
$E_{-6}=-\frac{3^{3/4}}{32b^{3/2}}\int\rho({\rm\bf r})\,\mathrm{d}^{3}{\rm\bf r}.$ (11)
Our first approximation assumes $\omega_{1}\approx\omega_{2}$, helping to simplify the function of Eq. (II) down to three variables
@@
and the new function of the squared distance
$v(s,\mu_{1},\mu_{2})=\frac{1}{(s+\mu_{1})(s+\mu_{2})\bigl(s+\tfrac{1}{2}(\mu_{1}+\mu_{2})\bigr)}.$ (15)
@@
The assumption $\omega_{1}\approx\omega_{2}$ can be justified by the distribution of $\omega({\rm\bf r})$ values of Eq. (6) in typical molecules that can be studied numerically, and in the same way we see a rather narrow distribution of $\mu({\rm\bf r})$ values of Eq. (14).
Our next approximation deals with the function (15)
@@
$\tilde{v}(s,\mu,\nu,\beta)=\frac{1}{(s+\beta)^{3}}+\sum_{m=4}^{n}\frac{c_{m}(\mu,\nu,\beta)}{(s+\beta)^{m}},$ (18)
where $\beta$ should be a system-independent parameter. At first we hoped the simplest one with only the first term might be ideal: it has the $r^{-6}$-tail the dispersion functionals are all about, and only one spatial distribution (13) is needed. This hope was forlorn after the tests on noble-gas homo- and heterodimers: the value of $\beta$ had to vary from system to system by more than twofold to reproduce well enough the potential curves. Thus the length-scale distribution (14) should be somehow accounted for.
@@
What also matters is the sum rule — also to leave the term (3) unchanged from (11). From the diagonal ($\mu=\nu$) case,
@@
For $\mu\neq\nu$ we take the weighted arithmetic and geometric mean for the last term,
@@
Even though $\mu({\rm\bf r})\approx\beta$ almost everywhere (as numerical studies show), it can be too big or too small somewhere, for greater numerical stability we prefer to use safer smoothed values
@@
$\theta(x,c,a)=c+ac\tanh\left(\frac{x-c}{ac}\right)$ (31)
(where $c$ sets the typical value and $a$ controls the spread), with an adjustable parameter $\alpha$ (see below).
(We are well aware that our approximation is (one of) the simplest, and more accurate ones can be developed applying the idea of Eq. (18) directly to $v_{0}$ of Eq. (II),
$\tilde{v}_{0}(s,\omega_{1},\kappa_{1},\omega_{2},\kappa_{2})=\sum_{m=3}^{n}\frac{c_{m}(\omega_{1},\kappa_{1},\omega_{2},\kappa_{2},\beta)}{(s+\beta)^{m}}.$ (32)
For example, using the approximation
$\frac{2}{p+q}\approx\frac{2}{\sqrt{pq}}-\frac{1}{2p}-\frac{1}{2q}$ (33)
we might get
$c_{3}=\frac{1}{\omega_{1}^{3/2}\omega_{2}^{3/2}}-\frac{1}{4\omega_{1}^{2}\omega_{2}}-\frac{1}{4\omega_{1}\omega_{2}^{2}},$ (34)
and so on, even with $n>5$ in Eq. (32), but we will not do it here.)
@@
are computed using the points $\{{\rm\bf r}_{k}\}$ of a three-dimensional integration grid with weights $\{w_{k}\}$, the fast multipole methods Rokhlin (1985) should help here.
@@ III Calculations
Homo- and heteroatomic noble-gas dimers are our favorite model system for testing dispersion functionals. We use our long-range-corrected version Laikov (2019c) of the PBE Perdew et al. (1996) exchange functional, our two-component scalar-relativistic approximation Laikov (2019a), and our L2a_3 basis set Laikov (2019b), and calculate the values shown in Table 1.
@@ IV Conclusions
A plane-wave implementation for solid-state systems should be straightforward using the known methods Sabatini et al. (2013).
V Data availability
The data that support the findings of this study are available within the article.